# Bayes Adaptive Monte Carlo Tree Search for Offline Model-based Reinforcement Learning

### v1

@@ Appendix A BAMCP
Appendix B Alternative Design Choices for Continuous BAMCP
@@
The terms labeled in blue in Algorithm 1 have alternative design choices. Empirical comparisons among these alternatives are reserved for future work.
@@
$\tilde{Q}((s,h),x)$ in the exploration strategy, i.e., $a\leftarrow\arg\max_{x\in C((s,h))}\tilde{Q}((s,h),x)$, could take various forms. For instance, in (Couëtoux et al. (2011); Guez et al. (2013); Lee et al. (2020)), $\tilde{Q}((s,h),x)=Q((s,h),x)+c\sqrt{\frac{\log N((s,h))}{N((s,h),x)}}$; in PUCT (Auger et al. (2013)), $\tilde{Q}((s,h),x)=Q((s,h),x)+\sqrt{\frac{N((s,h))^{e(d)}}{N((s,h),x)}}$, where $e(d)$ is a schedule of coefficients related to the search depth $d$; in Sampled MuZero (a variant of MuZero that can be applied in continuous action spaces (Hubert et al. (2021))), $\tilde{Q}((s,h),x)=Q((s,h),x)+\hat{\pi}(x|(s,h))\frac{\sqrt{N((s,h))}}{1+N((s,h),x)}\left(c_{1}+\log\left(\frac{N((s,h))+c_{2}+1}{c_{2}}\right)\right)$. Here, $c$, $c_{1}$, $c_{2}$ are hyperparameters, $\hat{\pi}=\hat{\beta}\pi^{1-1/\tau}$ is a sample policy defined upon the real policy $\pi$. In particular, at each decision point $(s,h)$, Sampled MuZero would sample $M$ actions $\{a_{i}\}$ from the distribution $\pi^{1/\tau}$ and accordingly define $\hat{\beta}(a|(s,h))=\sum_{i}\mathds{1}_{a_{i}=a}/M$, where $\tau>0$ is a temperature hyperparameter. Thus, Sampled MuZero does not adopt progressive widening like ours. Following BAMCP, we adopt the first definition of $\tilde{Q}((s,h),x)$, though it could potentially be improved with other choices. In addition, as an implementation trick (Hamrick et al. (2021)), the Q estimates are usually normalized into $\bar{Q}\in[0,1]$ before being used to calculate $\tilde{Q}$ as above. The normalized estimates can be computed as $\bar{Q}((s,h),x)=\frac{Q((s,h),x)-Q_{\text{min}}}{Q_{\text{max}}-Q_{\text{min}}}$, where $Q_{\text{max}}$ and $Q_{\text{min}}$ are the maximum and minimum Q values observed in the search tree so far.
@@
As the planning/search result, $\pi_{\text{ret}}$ can take multiple forms. In (Guez et al. (2013); Sunberg & Kochenderfer (2018); Lee et al. (2020)), $\pi_{\text{ret}}((s,h))=\arg\max_{a\in C((s,h))}Q((s,h),a)$; in (Sampled) MuZero, $\pi_{\text{ret}}(a|(s,h))=\frac{N((s,h),a)^{1/\tau}}{\sum_{x\in C((s,h))}N((s,h),x)^{1/\tau}}$; in ROSMO (a variant of MuZero with improved performance in offline scenarios (Liu et al. (2023))), $\pi_{\text{ret}}(a|(s,h))\propto\pi(a|(s,h))\exp(Q((s,h),a)-V((s,h)))$. Here, $\tau\in(0,1]$ is a temperature parameter and decays with the training process, ensuring the action selection becomes greedier. We select the second form for $\pi_{\text{ret}}$ in Algorithm 1. This is because (1) as described in PUCT, the returned action should be the most visited one, which is not necessarily the one with the highest Q value, and (2) ROSMO adopts one-step look-ahead rather than deep tree search at each root node, which does not align with our approach.
@@
As for the conditions of double progressive widening, PUCT designs $\alpha$ and $\beta$ to be functions of the search depth $d$, while UCT-DPW (Couëtoux et al. (2011)) utilizes a different set of conditions: $\left\lceil{K_{a}N((s,h))^{\alpha}}\right\rceil\geq|C((s,h))|$, $\left\lceil{K_{s}N((s,h),a)^{\beta}}\right\rceil\geq|C((s,h),a)|$, where $K_{a},K_{s},\alpha,\beta$ are all constant hyperparameters. When the progressive widening condition for sampling the next state is not satisfied, either the least visited node in $C((s,h),a)$ can be selected (following PUCT), or a random node can be sampled from $C((s,h),a)$ following a distribution proportional to the number of visits (following UCT-DPW). As shown in Algorithm 1, we follow the designs of PUCT, but keep $\alpha$ and $\beta$ as constants for simplicity in hyperparameter fine-tuning.
@@
Finally, the condition for continuing the simulation procedure, i.e., $N((s,h))>1$, could potentially be replaced with $N((s,h),a)>1$ or $N((s^{\prime},hars^{\prime}))>0$. These conditions indicate that the nodes $(s,h)$, $(s,h,a)$, and $(s^{\prime},hars^{\prime})$ have been visited before, respectively. At the end of the simulation procedure, we can either apply rollouts, i.e., simulating a single path until the end of an episode, to estimate the expected value for a leaf node $(s,h)$, or directly use $V((s,h))$ as the estimation. The former approach is widely used in online planning algorithms (Guez et al. (2013); Sunberg & Kochenderfer (2018); Lee et al. (2020)), while the latter is used in iterative frameworks like MuZero.
Appendix C Key Hyperparameter Setup

### v4

@@ B.3.1 Analysis of Standard Random Nodes ($w_{j},j>0$)
$+\underbrace{|\mathbb{E}[R_{w}(j)]-V^{*}_{p}(w)|}_{\text{(c) Bias / Approximation Error}}+\underbrace{|\frac{n(i_{0})}{n(w)}R_{w}(i_{0})|}_{\text{(d) ``Newest'' Child Error}}$
We now bound each term:
(a) Bounding the Weight Deviation Error
This term is bounded by leveraging the fact that the children of a random node are explored in a nearly uniform manner (a consequence of the DPW mechanism, as shown in Lemma 1 of Auger et al. (2013). Following their derivation, this term is bounded by $O(n(w)^{-(1-\alpha)}+n(w)^{-\alpha})$.
(b) Bounding the Sampling Error of “Mature” Returns
This term represents the concentration error for the returns of the mature child nodes. The returns $R_{w}(i)$ are i.i.d. bounded random variables. Therefore, we can apply Hoeffding’s inequality. To align with the proof by Auger et al. (2013) and derive the resulting convergence rate for this layer, we select the deviation tolerance $t$ as:
$t:=n(w)^{-\frac{\gamma_{d}}{1+3\gamma_{d}}}$
With this choice, the application of Hoeffding’s inequality shows that the error for this term is bounded by $O(n(w)^{-\frac{\gamma_{d}}{1+3\gamma_{d}}})$ and that this bound holds exponentially surely.
(c) Bounding the Bias / Approximation Error
This term is unique to our proof and requires our induction hypothesis.
$|\mathbb{E}[R_{w}(j)]-V^{*}_{p}(w)|$ $=|\mathbb{E}[r_{j}+\gamma\hat{V}(j)]-\mathbb{E}[r_{j}+\gamma V^{*}_{p}(j)]|$
$=\gamma|\mathbb{E}[\hat{V}(j)-V^{*}_{p}(j)]|$
$\leq\gamma\mathbb{E}[|\hat{V}(j)-V^{*}_{p}(j)|]$
By the induction hypothesis for nodes at depth $d$, we have $|\hat{V}(j)-V^{*}_{p}(j)|\leq C_{d}n(j)^{-\gamma_{d}}+f_{d}(\epsilon_{V})$. Substituting this into the inequality gives:
$|\mathbb{E}[R_{w}(j)]-V^{*}_{p}(w)|\leq\gamma\mathbb{E}[C_{d}n(j)^{-\gamma_{d}}]+\gamma f_{d}(\epsilon_{V})\text{ e.s. in }n(w)$
The first part of this bound is a decaying term, while the second part shows how the bias propagates.
(d) Bounding the “Newest” Child Error
Following the same logic as Auger et al. (2013), this term is bounded by $O(n(w)^{-\alpha})$ because the newest child has been visited only a small number of times.
Conclusion for Standard Nodes
Combining these bounds, the overall convergence rate is determined by the slowest-converging term. Following the derivation in the PUCT paper, setting the exploration parameter $\alpha=\frac{3\gamma_{d}}{1+3\gamma_{d}}$, establishes the slowest convergence rate as $\frac{\gamma_{d}}{1+3\gamma_{d}}$. The bias term is propagated from the layer below. Thus, for any standard random node, the Cons($\gamma,f,d$) property holds.
B.3.2 Analysis of the First-Child Random Node ($w_{0}$)
For the first child $w_{0}$, the value estimate includes a special term from its initial visit. On the first visit, the algorithm samples a next state, which we denote $s_{0}^{\prime}$, and receives a corresponding reward $r_{0}$. The next state’s value is thus initially estimated using the learned function, $V(s_{0}^{\prime})$. The error analysis for this random node, $|\hat{V}(w_{0})-V^{*}_{p}(w_{0})|$, will contain the same four error terms as in the standard case (for the subsequent backups), plus an additional error term originating from the first visit:
$\text{Additional Error}=\left|\frac{1}{n(w_{0})}(r_{0}+\gamma V(s^{\prime}_{0}))-\frac{1}{n(w_{0})}V^{*}_{p}(w_{0})\right|$
Since $r_{0}$, $V(s^{\prime}_{0})$, and $V^{*}_{p}(w_{0})$ are all bounded constants, this additional error term is bounded by $O(n(w_{0})^{-1})$.
The overall convergence rate for $w_{0}$ is determined by the minimum of the rates of all error components. From the PUCT derivation, we know that $\gamma_{d}\leq 1$, which implies $\gamma_{d-1/2}=\frac{\gamma_{d}}{1+3\gamma_{d}}<1$. Therefore, the convergence rate of the additional error term, $O(n(w_{0})^{-1})$, is faster than the bottleneck rate of the standard MCTS error terms, $O(n(w_{0})^{-\gamma_{d-1/2}})$. This special term does not change the overall convergence rate. Thus, the Cons($\gamma,f,d$) property also holds for the first-child random node $w_{0}$ with the exact same parameters.
B.3.3 Conclusion for this Section
We have shown that for any random node $w$ at depth $d-\frac{1}{2}$, if its children at depth $d$ satisfy the Cons($\gamma,f,d$) property, then $w$ itself also satisfies the property. The parameters of the property propagate upward according to the following recursive formulas:
  • $\gamma_{d-\frac{1}{2}}=\frac{\gamma_{d}}{1+3\gamma_{d}}$
  • $f_{d-\frac{1}{2}}(\epsilon_{V})=\gamma f_{d}(\epsilon_{V})$
B.4 Inductive Step: From Random to Decision Nodes
In this section, we perform the second inductive step. We assume the induction property Cons($\gamma,f,d$) holds for all random nodes $w$ at depth $d+\frac{1}{2}$. Our objective is to prove that the property also holds for their parent nodes, which are the decision nodes $z$ at depth $d$. The proof adapts the structure of Section 6 in Auger et al. (2013), carefully accounting for the bias term $f_{d+\frac{1}{2}}(\epsilon_{V})$ from our induction hypothesis. First we establish an upper bound on $\hat{V}(z)-V^{*}_{p}(z)$, and then a lower bound.
B.4.1 Upper Bound on the Error
To establish an upper bound on $\hat{V}(z)-V^{*}_{p}(z)$, we partition the children $w_{i}$ of $z$ into two classes, with a parameter $\epsilon<1-\alpha$ to be determined later:
  • Class I: Children $w_{i}$ with few visits, $n(w_{i})\leq n(z)^{1-\alpha-\epsilon}$.
  • Class II: All other children, for which $n(w_{i})>n(z)^{1-\alpha-\epsilon}$.
For Class I, the number of children is at most $n(z)^{\alpha}$, so their total contribution is bounded by $\frac{n(z)^{\alpha}\cdot n(z)^{1-\alpha-\epsilon}}{n(z)}=n(z)^{-\epsilon}$. For Class II, we apply the induction hypothesis for nodes at depth $d+\frac{1}{2}$:
$\sum_{i\in\text{Class II}}\frac{n(w_{i})}{n(z)}(\hat{V}(w_{i})-V^{*}_{p}(z))$ $\leq\sum_{i\in\text{Class II}}\frac{n(w_{i})}{n(z)}(\hat{V}(w_{i})-V^{*}_{p}(w_{i}))$
$\leq\sum_{i\in\text{Class II}}\frac{n(w_{i})}{n(z)}(C_{d+\frac{1}{2}}n(w_{i})^{-\gamma_{d+\frac{1}{2}}}+f_{d+\frac{1}{2}}(\epsilon_{V}))$
$\leq C_{d+\frac{1}{2}}\left(n(z)^{1-\alpha-\epsilon}\right)^{-\gamma_{d+\frac{1}{2}}}+f_{d+\frac{1}{2}}(\epsilon_{V})$
Combining the two parts, we get:
$\hat{V}(z)-V^{*}_{p}(z)\leq n(z)^{-\epsilon}+C_{d+\frac{1}{2}}n(z)^{-\gamma_{d+\frac{1}{2}}(1-\alpha-\epsilon)}+f_{d+\frac{1}{2}}(\epsilon_{V})$
To optimize this bound, we balance the exponents by choosing $\epsilon=\frac{\gamma_{d+\frac{1}{2}}(1-\alpha)}{1+\gamma_{d+\frac{1}{2}}}$. This leads to the final upper bound for the error:
$\hat{V}(z)-V^{*}_{p}(z)\leq(1+C_{d+\frac{1}{2}})n(z)^{-\frac{\gamma_{d+\frac{1}{2}}(1-\alpha)}{1+\gamma_{d+\frac{1}{2}}}}+f_{d+\frac{1}{2}}(\epsilon_{V})$
B.4.2 Lower Bound on the Error
The proof for the lower bound is more involved. It relies on the parameter definitions from the PUCT paper, as well as two key results regarding the UCT exploration mechanism, which we state here for completeness before proceeding.
Lemma B.5 (Adapted from Lemma 3, Auger et al. (2013)).
Consider a bandit setting where the score of a child node $w_{i}$ at time $n$ is computed by the UCT rule:
$sc_{n}(w_{i})=\hat{V}_{n}(w_{i})+\sqrt{\frac{g(n)}{n(w_{i})}}$
where $g(n)$ is a non-decreasing exploration function. If $i$ denotes the $i^{th}$ constucted child, for all $n>i^{\frac{1}{\alpha(1-\alpha)}}$ we have
$n(w_{i})\geq\frac{1}{4}\min\left(g(n^{1-\alpha}),n^{1-\alpha}\right).$
Corollary B.6 (Adapted from Corollary 3, Auger et al. (2013)).
For the exploration function $g(n)=n^{e}$ with $0<e<1$, we obtain
$n(w_{i})\geq\frac{1}{4}n^{e(1-\alpha)}\text{ if }i\leq n^{\alpha(1-\alpha)}.$ (8)
With these results, we can define the parameters as functions of $\gamma_{d+\frac{1}{2}}$ and $p$ (from Assumption B.2) for this layer and proceed with the three-step proof structure.
  • Progressive widening coefficient: $\alpha:=\frac{\gamma_{d+\frac{1}{2}}}{1+4\gamma_{d+\frac{1}{2}}}$
  • Exploration coefficient: $e:=\frac{1}{2p\left(1+4\gamma_{d+\frac{1}{2}}\right)}$
  • Auxiliary proof parameters: $\xi:=\frac{1}{1+e\gamma_{d+\frac{1}{2}}(1-\alpha)}$ and $\Delta:=\left(\frac{1}{4}n(z)^{\xi e(1-\alpha)}\right)^{-\gamma_{d+\frac{1}{2}}}$
Step 1: Existence of a near-optimal child
Based on Assumption B.2 (Policy Regularity), a direct adaptation of the proof in Auger et al. (2013) shows that a near-optimal child node is discovered with high probability. Formally: Exponentially surely in $n(z)$, there exists at time $\lceil n(z)^{\xi(1-\alpha)}\rceil$ a child $w_{0}$ of $z$ such that
$V^{*}_{p}(w_{0})\geq V^{*}_{p}(z)-\Delta\text{ and }i_{0}\leq n(z)^{\xi(1-\alpha)\alpha}$ (9)
Step 2: Score of children selected at later stages
The objective of this step is to show that any child selected after time $n(z)^{\xi}$ must have a high score. Let $n^{\prime}$ be such that $n^{\xi}\leq n^{\prime}\leq n$. And, by Corollary B.6,
$n^{\prime}(w_{0})\geq\frac{1}{4}n^{\prime e(1-\alpha)}\geq\frac{1}{4}n^{\xi e(1-\alpha)}.$
First, according to Corollary B.6, the near-optimal child $w_{0}$ discovered in Step 1 will be visited a number of times $n(w_{0})$ that grows polynomially with $n(z)$.
We can therefore apply our induction hypothesis Cons($\gamma,f,d$) to $w_{0}$, with the Eq.(9), there exists a $C^{\prime}>0$ that:
$\hat{V}(w_{0})$ $\geq V^{*}_{p}(w_{0})-C^{\prime}\left(\frac{1}{4}n^{\xi e(1-\alpha)}\right)^{-\gamma_{d+\frac{1}{2}}}-f_{d+\frac{1}{2}}(\epsilon_{V})$
$\geq V^{*}_{p}(z)-(1+C^{\prime})\Delta-f_{d+\frac{1}{2}}(\epsilon_{V})$
Now, consider any child $w_{1}$ selected by the UCT rule at a later time $n^{\prime}\geq n(z)^{\xi}$. Its score must be at least as high as the score of $w_{0}$. This gives the key inequality:
$\hat{V}(w_{1})+\sqrt{\frac{n^{\prime e}}{n^{\prime}(w_{1})}}\geq\hat{V}(w_{0})+\sqrt{\frac{n^{\prime e}}{n^{\prime}(w_{0})}}$
Combining these results, we can establish a lower bound on the estimated value $\hat{V}(w_{1})$ of any child selected at a late stage of the search.
$\hat{V}(w_{1})+\sqrt{\frac{n}{n^{\prime}(w_{1})}}\geq V^{*}_{p}(z)-(1+C^{\prime})\Delta-f_{d+\frac{1}{2}}(\epsilon_{V})$
As shown by Auger et al. (2013), this property holds exponentially surely in $n(z)$.
Step 3: Bounding the value estimate $\hat{V}(z)$
Consider a child $w_{1}$ selected after $n(z)^{\xi}$. By the previous step, exponentially in $n(z)$, this child must either satisfy
$\sqrt{\frac{n^{e}}{n(w_{1})}}\geq\Delta$ (10)
$\text{or }\hat{V}(w_{1})\geq V(z)-(2+C^{\prime})\Delta-f_{d+\frac{1}{2}}(\epsilon_{V})$ (11)
We now bound the average value $\hat{V}(z)$ by partitioning the children of $z$ into three categories:
  1. 1.
    Children visited only before time $n(z)^{\xi}$.
  2. 2.
    Children visited after $n(z)^{\xi}$ satisfying (10)
  3. 3.
    Children visited after $n(z)^{\xi}$ satisfying (11).
We bound the contribution of each category to the total error sum. The contribution from Category 1 is bounded by $O(n(z)^{\xi-1})$, since we have
$\left|\sum_{w\text{ in cat.1}}\frac{n(w)}{n(z)}\left(\hat{V}(w)-V^{*}_{p}(z)\right)\right|\leq\frac{\sum_{w\text{ in cat.1}}n(w)}{n(z)}\leq\frac{n(z)^{\xi}}{n(z)}$
For Category 2, the definition of the category implies $n(w_{i})\leq n(z)^{e}/\Delta^{2}$. Since there are at most $n(z)^{\alpha}$ children, their total contribution to the error sum is bounded by $O(n(z)^{\alpha+e-1}/\Delta^{2})$. For Category 3, the bound on $\hat{V}(w_{i})$ from Step 2 introduces the crucial $f_{d+1/2}(\epsilon_{V})$ term.
Combining the bounds from all three categories gives the final lower bound on the error. Following the argument in Auger et al. (2013):
$\hat{V}(z)-V^{*}_{p}(z)\geq\underbrace{-(2+C^{\prime})\Delta(1-n(z)^{\xi-1})-f_{d+\frac{1}{2}}(\epsilon_{V})}_{\text{cat.3}}-\underbrace{n(z)^{\xi-1}}_{\text{cat.1}}-\underbrace{\frac{n(z)^{\alpha+e-1}}{\Delta^{2}}}_{\text{cat.2}}$
By comparing the magnitudes of the decaying terms with the chosen parameter definitions, as is done in Auger et al. (2013), this complex expression simplifies to the desired form.
B.4.3 Conclusion for this Section
By combining the results from the upper and lower bound analyses, we conclude that the total error at a decision node $z$ is bounded as:
$|\hat{V}(z)-V^{*}_{p}(z)|\leq C_{d}n(z)^{-\gamma_{d}}+f_{d+\frac{1}{2}}(\epsilon_{V})$
This demonstrates that the induction property Cons($\gamma,f,d$) holds for depth $d$. The parameters propagate upward according to the following recursive formulas, which are identical to those derived by Auger et al. (2013):
  • $C_{d}=(5+C^{\prime})4^{\gamma_{d+\frac{1}{2}}}$
  • $\gamma_{d}=\frac{\gamma_{d+\frac{1}{2}}}{1+7\gamma_{d+\frac{1}{2}}}$
  • $f_{d}(\epsilon_{V})=f_{d+\frac{1}{2}}(\epsilon_{V})$
B.5 Conclusion of the proof
In the preceding sections, we have established all the necessary components for our backward induction argument.
  • In Section B.2, we established the base case, proving that the induction property Cons($\gamma,f,d$) holds for random nodes at depth $d=d_{max}-\frac{1}{2}$, with initial parameters $\gamma_{d_{max}-\frac{1}{2}}=\frac{1}{3}$ and $f_{d_{max}-\frac{1}{2}}(\epsilon_{V})=\epsilon_{V}$.
  • In Section B.3, we proved the inductive step from decision nodes up to random nodes, deriving the recursions $\gamma_{d-\frac{1}{2}}=\frac{\gamma_{d}}{1+3\gamma_{d}}$ and $f_{d-\frac{1}{2}}(\epsilon_{V})=\gamma f_{d}(\epsilon_{V})$.
  • In Section B.4, we proved the inductive step from random nodes up to decision nodes, deriving the recursions $\gamma_{d}=\frac{\gamma_{d+\frac{1}{2}}}{1+7\gamma_{d+\frac{1}{2}}}$ and $f_{d}(\epsilon_{V})=f_{d+\frac{1}{2}}(\epsilon_{V})$.
With the base case proven and both inductive steps validated, we conclude by backward induction that the property Cons($\gamma,f,d$) holds for all depths up to the root of the search tree ($d=0$).
By solving the recursion for the bias term, we can find the accumulated error at the root. The base error at depth $d_{max}-\frac{1}{2}$ is $f_{d_{max}-\frac{1}{2}}(\epsilon_{V})=\epsilon_{V}$. The error term is multiplied by a discount factor $\gamma$ each time it propagates from a decision node layer up to a random node layer. This occurs $d_{max}$ times from the leaves to the root. Therefore, the bias term at the root node ($d=0$) is $f_{0}(\epsilon_{V})=\gamma^{d_{max}}\epsilon_{V}$.
Specifically, for the root node $z_{0}$, its total error is bounded by:
$|\hat{V}(z_{0})-V^{*}_{p}(z_{0})|\leq C_{0}n(z_{0})^{-\gamma_{0}}+\gamma^{d_{max}}\epsilon_{V}$
where $\gamma_{0}>0$ and $C_{0}$ are constants derived from the recursions. This shows that as the number of simulations $E=n(z_{0})\rightarrow\infty$, the estimated value of the root node converges to a ball of radius $\gamma^{d_{max}}\epsilon_{V}$ around the true optimal value of the P-MDP. Notably, the size of this error ball, which stems from our function approximator, shrinks exponentially with the search depth $d_{max}$.
This completes the proof of Theorem 4.1, establishing that Continuous BAMCP is a consistent planner for the P-MDP. ∎
Appendix C Alternative Design Choices for Continuous BAMCP
@@
The terms labeled in blue in Algorithm 1 have alternative design choices, which are introduced as below. Empirical comparisons among these alternatives are reserved for future work.
@@
$\tilde{Q}((s,h),x)$ in the exploration strategy, i.e., $a\leftarrow\arg\max_{x\in C((s,h))}\tilde{Q}((s,h),x)$, could take various forms. For instance, in Couëtoux et al. (2011); Guez et al. (2013); Lee et al. (2020), $\tilde{Q}((s,h),x)=Q((s,h),x)+c\sqrt{\frac{\log N((s,h))}{N((s,h),x)}}$; in PUCT (Auger et al., 2013), $\tilde{Q}((s,h),x)=Q((s,h),x)+\sqrt{\frac{N((s,h))^{e(d)}}{N((s,h),x)}}$, where $e(d)$ is a schedule of coefficients related to the search depth $d$; in Sampled MuZero (a variant of MuZero that can be applied in continuous action spaces (Hubert et al., 2021)), $\tilde{Q}((s,h),x)=Q((s,h),x)+\hat{\pi}(x|(s,h))\frac{\sqrt{N((s,h))}}{1+N((s,h),x)}\left(c_{1}+\log\left(\frac{N((s,h))+c_{2}+1}{c_{2}}\right)\right)$. Here, $c$, $c_{1}$, $c_{2}$ are hyperparameters, $\hat{\pi}=\hat{\beta}\pi^{1-1/\tau}$ is a sample policy defined upon the real policy $\pi$. In particular, at each decision point $(s,h)$, Sampled MuZero would sample $M$ actions $\{a_{i}\}$ from the distribution $\pi^{1/\tau}$ and accordingly define $\hat{\beta}(a|(s,h))=\sum_{i}\mathds{1}_{a_{i}=a}/M$, where $\tau>0$ is a temperature hyperparameter. Thus, Sampled MuZero does not adopt progressive widening like ours. Following BAMCP, we adopt the first definition of $\tilde{Q}((s,h),x)$, though it could potentially be improved with other choices. In addition, as an implementation trick (Hamrick et al., 2021), the Q estimates are usually normalized into $\bar{Q}\in[0,1]$ before being used to calculate $\tilde{Q}$ as above. The normalized estimates can be computed as $\bar{Q}((s,h),x)=\frac{Q((s,h),x)-Q_{\text{min}}}{Q_{\text{max}}-Q_{\text{min}}}$, where $Q_{\text{max}}$ and $Q_{\text{min}}$ are the maximum and minimum Q values observed in the search tree so far.
@@
As the planning/search result, $\pi_{\text{ret}}$ can take multiple forms. In Guez et al. (2013); Sunberg and Kochenderfer (2018); Lee et al. (2020), $\pi_{\text{ret}}((s,h))=\arg\max_{a\in C((s,h))}Q((s,h),a)$; in (Sampled) MuZero, $\pi_{\text{ret}}(a|(s,h))=\frac{N((s,h),a)^{1/\tau}}{\sum_{x\in C((s,h))}N((s,h),x)^{1/\tau}}$; in ROSMO (a variant of MuZero with improved performance in offline scenarios (Liu et al., 2023)), $\pi_{\text{ret}}(a|(s,h))\propto\pi(a|(s,h))\exp(Q((s,h),a)-V((s,h)))$. Here, $\tau\in(0,1]$ is a temperature parameter and decays with the training process, ensuring the action selection becomes greedier. We select the second form for $\pi_{\text{ret}}$ in Algorithm 1. This is because (1) as described in PUCT, the returned action should be the most visited one, which is not necessarily the one with the highest Q value, and (2) ROSMO adopts one-step look-ahead rather than deep tree search at each root node, which does not align with our approach.
@@
As for the conditions of double progressive widening, PUCT designs $\alpha$ and $\beta$ to be functions of the search depth $d$, while UCT-DPW (Couëtoux et al., 2011) utilizes a different set of conditions: $\left\lceil{K_{a}N((s,h))^{\alpha}}\right\rceil\geq|C((s,h))|$, $\left\lceil{K_{s}N((s,h),a)^{\beta}}\right\rceil\geq|C((s,h),a)|$, where $K_{a},K_{s},\alpha,\beta$ are all constant hyperparameters. Moreover, when the progressive widening condition for sampling the next state is not satisfied, either the least visited node in $C((s,h),a)$ can be selected (following PUCT), or a random node can be sampled from $C((s,h),a)$ according to a distribution proportional to the number of visits (following UCT-DPW). As shown in Algorithm 1, we follow the designs of PUCT, but keep $\alpha$ and $\beta$ as constants for simplicity in hyperparameter fine-tuning.
@@
Finally, the condition for continuing the simulation procedure, i.e., $N((s,h))>1$, could potentially be replaced with $N((s,h),a)>1$ or $N((s^{\prime},hars^{\prime}))>0$. These conditions indicate that the nodes $(s,h)$, $(s,h,a)$, and $(s^{\prime},hars^{\prime})$ have been visited before, respectively. At the end of the simulation procedure, we can either apply rollouts, i.e., simulating a single path until the end of an episode, to estimate the expected value for a leaf node $(s,h)$, or directly use $V((s,h))$ as the estimation. The former approach is widely used in online planning algorithms (Guez et al., 2013; Sunberg and Kochenderfer, 2018; Lee et al., 2020), while the latter is used in policy iteration frameworks like ours.
Appendix D Key Hyperparameter Setup